# Choose Your Model Size: Any Compression by a Single Gradient Descent

### v1

@@ Appendix D Further Experiments and Ablations
In this section, we present several supplementary experiments analyzing the impact of some key algorithmic components and design choices of ACIP in more detail.
D.1 Impact of Low-Rank Adapters
The primary purpose of the low-rank adapters used in ACIP is to correct compression errors on-the-fly during the optimization. A surprising finding of our work is that the final adapters are “universal” in the sense that they can be used across all seen compression levels. While we expect that other PEFT-style approaches would lead to similar findings, it is natural to ask how ACIP would perform without any correction, i.e., just the mask parameters are tuned according to (8). This ablation study is shown in Figure A9. While performing significantly worse than with LoRA, we observe that the perplexity does not blow up and the results are even slightly better than SVD-LLM (see Table 1). This stable behavior of ACIP is closely related to our parameterization of the mask in (7) which ensures that the forward pass corresponds to the actual outputs of the pruned model with binary masks. On the other hand, the straight-through estimator still enables backpropagation.
@@ D.1 Impact of Low-Rank Adapters
D.2 Impact of the Stopping Criterion
In most experiments, we have used a compression ratio $r_{\text{stop}}$ of $0.4$ as stopping criterion, i.e., the pruning of masks is stopped if the size of the model is only 40% of the original one (measured in number of parameters of all target weight matrices). We have observed that at this point, the model performance has typically dropped so much that even a fine-tuned model would be of limited practical use.
Nevertheless, it is interesting to explore the sensitivity of compression-performance curves against different stopping ratios. The comparison shown in Figure A10 provides several insights in this respect: (i) “Forecasting” compressed models beyond the stopping ratio does not work very well, especially when stopping very early ($>0.8$). (ii) The predictive capacity of ACIP remains valid for even stronger stopping compression ratios than $0.4$. However, finding the largest reasonable stopping ratio is highly model-dependent. For less compressible models like LLaMA-3.1-8B, it could make sense to stop even earlier than $0.4$ (cf. Figure 1). In general, we hypothesize that older models are more compressible than new ones, as the latter “carry” more information per weight due to significantly more training data (Allen-Zhu & Li, 2024).
@@ D.2 Impact of the Stopping Criterion
D.3 Impact of the Score Map – Forecasting Pruning Patterns
Here, we pick up the observation from Section D.2 that forecasting the performance of compressed models beyond the stop ratio leads to inaccurate predictions, i.e., the model is compressed more strongly than it has been done by ACIP itself. However, it turns out that the score map itself exhibits a certain forecasting capability. To this end, we run ACIP as usual until a stop ratio is reached, say $r_{\text{stop}}=0.4$, but we stop updating the score map earlier in the optimization process. A few compression-performance curves with this modification are reported in Figure A11. We observe very similar curve shapes even if the score map is frozen after only a tiny fraction of mask parameters was pruned. This underpins our intuition from Section 2.2.2 that the pruning path of each parameter is fully determined at very early stage of ACIP.
@@ D.3 Impact of the Score Map – Forecasting Pruning Patterns
D.4 Impact of the Score Map – A Trivial One Does Not Work
There are certainly alternative ways to design useful score maps. For example, simply accumulating the gradients of all mask parameters entrywise over an ACIP-run works equally well as the strategy proposed in Section 2. It is therefore valid to ask whether one could even design score maps without any optimization. We demonstrate that perhaps the most obvious approach, namely setting the score map equal to the singular values of the weight matrices, does not work very well. Figure A12 shows that this training-free approach does not produce any reasonable compressed models and decent performance cannot be easily recovered with LoRA-finetuning. This simple experiment confirms that designing useful score maps is not a trivial endeavour and requires a carefully crafted algorithmic approach.
@@ D.4 Impact of the Score Map – A Trivial One Does Not Work
D.5 Impact of Post-Tuning
Our main experiments are performed with 1K post-tuning steps in ACIP (see the description in Section 2.2.2). Figure A13 shows analogous compression-performance trade-off curves for fewer or no post-tuning steps. We observe that post-tuning can indeed notably increase performance for higher compression ratios.
@@ D.5 Impact of Post-Tuning
D.6 Impact of Individual Layers – Example of LLaMA2-13B
As pointed out in the caption of Table A2, the linear layers targeted by ACIP were slightly modified for LLaMA2-13, namely all up projection layers were ignored. Figure A14 shows what would happen if they are compressed as well. While the performance predictions for $\geq 0.6$ look decent, the perplexity explodes for stronger compression; note that even additional fine-tuning does not recover a reasonable performance in this situation. We hypothesize that ACIP has pruned one or more singular values of the up projection layers that are crucial for model’s integrity. This finding might be related to the recent work by Yu et al. (2024) on pruning so-called super weights. In any case, ACIP is capable of revealing this undesirable behavior as demonstrated in Figure A14.
@@ D.6 Impact of Individual Layers – Example of LLaMA2-13B
D.7 Examples of Score Maps Generated by ACIP
Figure A15 and Figure A16 show two typical score maps generated by ACIP for LLaMA-7B and Qwen2.5-7B, respectively. A characteristic feature is that attention layers can be pruned more aggressively than the MLP layers. Similarly, we observe non-uniform pruning patterns for layers of the same type across all transformer layers. This confirms the findings of (Yuan et al., 2024; Jaiswal et al., 2024) and demonstrates that non-uniform structured compression can be achieved without any feature engineering.
@@ D.7 Examples of Score Maps Generated by ACIP
Figure A15: Example score maps generated by ACIP for LLaMA-7B. The negative values (cf. Algorithm 1) are normalized to $-1$ for the purpose of visualization.
@@
Figure A16: Example score maps generated by ACIP for Qwen2.5-7B. The negative values (cf. Algorithm 1) are normalized to $-1$ for the purpose of visualization.
D.8 Examples of Generated Text by Compressed Models
Table A4 shows examples of generated text by compressed versions of LLaMA-7B.
@@ D.8 Examples of Generated Text by Compressed Models
70% Apr 04, 2018 at 12:05 pm In the last 5-6 years, Natural Language Systems have been an important direction for my work and the research I do in the course of my PhD. I have been using them to teach myself Swedish and improve my Swedish by machine time in this post. Let’s see if I’ve got it! … Please tell me where is the concept of time in this poem. What is the significance of the poem ”I Am Wine” by H.W.C. Carson”? What does he mean by the poem title? What does the stairs, the wine, & the man have in common? What does he mean by ”I am wine”? What do you mean by ”I am wine”? …
@@
50% Jeffrey Huan, Shu-Hua Ji, Xiao-Hong Li, Xiao-Xiao Li, Layin Ma”s. 2019:5, Springer.0000.0001/13/8/19/12/720-22-F43. Miller A. ”The Structure of the Language in Unsupervised Deep-Processing” in Journal of the Acoustical Actors (v. 11(6). 2006:, pp. 278-295). Dong Ming, Sun Jin, He Mei-Chen,’ … Getting to this answer isn’t that hard. The concept of time is all about the study of time. The study of time isn’t so hard to get a hold on. When you’re younger, you can begin thinking of how much you could have had if you had known how much time you had…if only we had it back then …

### v2

@@ Appendix D Further Experiments and Ablations
In this section, we present several supplementary experiments analyzing the impact of some key algorithmic components and design choices of ACIP in more detail. Note that the most detailed analyses and ablations are carried out with LLaMA-7B as it was most extensively studied in previous research on structured weight pruning.
D.1 Efficiency Analysis
Table˜A5 reports several statistics on the efficiency of the ACIP algorithm and inference speed of compressed models. While these preliminary results do not immediately indicate gains in inference speed, we expect that further optimization like merging the low-rank adapters can compensate for the matrix-factorization overhead (one additional matrix-vector multiplication) and outperform the base model. Moreover, we note that compared to performance-size trade-offs, which are our main concern, analyzing inference speed-ups requires a very careful consideration about the hardware in use (accelerator model, parallel processing units, etc.) and measurement setup (sequence length, batch size, etc.).
Table A5: Efficiency analysis of ACIP for LLaMA-7B. The first three rows report the runtime and memory statistics of ACIP’s key steps (see Section˜2.2 and Figure˜2) both in terms of numbers and their qualitative asymptotics. Here, the model sizes are measured as (uncompressed) checkpoint sizes. “Runtime pruning” refers to the process of pruning the mask parameters to a desired compression ratio (revertible), whereas “Runtime compress” refers to the process of discarding pruned singular vectors and possibly unparametrizing linear layers, so that the model gets actually compressed (see Step 3 in Section˜2.2.3). The statistics of inference speed were obtained by generating new text of sequence length $64$ and batch size $64$. To measure FLOPs, we use the fvcore package and an input sequence of length $512$.
Stage Metric LLaMA-7B
ACIP Step 1 (Model Reparametrization) $\mathcal{O}(\text{\#Layers $\times$ SVD of Layer})$ Runtime [min] 4.95
Size parametrized model [GB] 19.71
Size base model [GB] 12.70
ACIP Step 2 (Scoring by Iterative Pruning) $\mathcal{O}(\text{\#Steps of Masks \& LoRA Updates})$ Runtime [min] 23.12
Reserved GPU memory peak [GB] 62.45
Steps / s 1.68
ACIP Step 3 (Any Compression) $\mathcal{O}(\text{\#Layers $\times$ Layer Input Dimension})$ Runtime pruning [s] 0.49
Runtime compress [s] 0.18
Inference at 40% Size Size model [GB] 5.47
Reserved GPU memory peak [GB] 25.68
Latency [s] 2.57
Tokens / s 1594.99
GigaFLOPs 1335.85
Inference at 70% Size Size model [GB] 9.10
Reserved GPU memory peak [GB] 29.43
Latency [s] 2.47
Tokens / s 1658.63
GigaFLOPs 2265.03
Inference at 100% Size (Original) Size model [GB] 12.70
Reserved GPU memory peak [GB] 32.79
Latency [s] 1.67
Tokens / s 2447.75
GigaFLOPs 3188.63
D.2 Impact of Low-Rank Adapters
The primary purpose of the low-rank adapters used in ACIP is to correct compression errors on-the-fly during the optimization. A surprising finding of our work is that the final adapters are “universal” in the sense that they can be used across all seen compression levels. While we expect that other PEFT-style approaches would lead to similar findings, it is natural to ask how ACIP would perform without any correction, i.e., just the mask parameters are tuned according to (8). This ablation study is shown in Figure˜A9. While performing significantly worse than with LoRA, we observe that the perplexity does not blow up and the results are even slightly better than SVD-LLM (see Table˜1). This stable behavior of ACIP is closely related to our parameterization of the mask in (7) which ensures that the forward pass corresponds to the actual outputs of the pruned model with binary masks. On the other hand, the straight-through estimator still enables backpropagation.
@@ D.2 Impact of Low-Rank Adapters
D.3 Impact of the Compression Rule
The Any Compression stage of ACIP (Step 3 in Section˜2.2.3) relies on a simple algorithmic rule: Prune as many singular values according to the learned score map $\boldsymbol{\rho}$ as needed for a desired compression rate $r$. Here, the tuned low-rank adapters $\boldsymbol{\Delta}$ are used across all compression levels, even for $r=1.0$, i.e., the size of the (uncompressed) base model. While the usage of low-rank adapters is helpful for simultaneous error correction along with iterative pruning (Step 2 in Section˜2.2.2), it can lead to a slight drop in performance for lower compression levels. This became already apparent in the ablation of Figure˜A9 above, where LoRA was fully omitted.
Figure A10: Compression-performance trade-off curves for LLaMA-7B on C4 with and without resetting linear layers if they are incompressible for a given target compression rate $r$.
Figure˜A10 shows that this performance gap can be reduced by a refinement of Step 3 in ACIP: If a linear layer $l$ is not compressible for a given rate $r$ (i.e., a low-rank factorization would not save any parameters), we reset the corresponding mask parameters $\mathbf{p}_{l}$ to $1.0$ and disable the low-rank adapter $\boldsymbol{\Delta}_{l}$. In this way, the layer is fully reset, and for $r=1.0$, we exactly recover the base model. Note that this adapted rule is fully reversible and therefore still allows for Any Compression.
D.4 Impact of the Stopping Criterion
In most experiments, we have used $r_{\text{stop}}=0.4$ as maximum reasonable compression ratio, i.e., the pruning of masks is stopped if the size of the model is only 40% of the original one (measured in number of parameters of all target weight matrices). We have observed that at this point, the model performance has typically dropped so much that even a fine-tuned model would be of limited practical use.
Nevertheless, it is interesting to explore the sensitivity of compression-performance curves against different stopping ratios. The comparison shown in Figure˜A11 provides several insights in this respect: (i) “Forecasting” compressed models beyond the stopping ratio does not work very well, especially when stopping very early (>0.8>0.8). (ii) The predictive capacity of ACIP remains valid for even stronger stopping compression ratios than $0.4$. However, finding the largest reasonable stopping ratio is highly model-dependent. For less compressible models like LLaMA-3.1-8B, it could make sense to stop even earlier than $0.4$ (cf. Figure˜4). In general, we hypothesize that older models are more compressible than new ones, as the latter “carry” more information per weight due to significantly more training data Allen-Zhu & Li (2024).
@@ D.4 Impact of the Stopping Criterion
D.5 Impact of the Score Map – Forecasting Pruning Patterns
@@
Here, we pick up the observation from Section˜D.4 that forecasting the performance of compressed models beyond the stop ratio leads to inaccurate predictions, i.e., the model is compressed more strongly than it has been done by ACIP itself. However, it turns out that the score map itself exhibits a certain forecasting capability. To this end, we run ACIP as usual until a stop ratio is reached, say $r_{\text{stop}}=0.4$, but we stop updating the score map earlier in the optimization process. A few compression-performance curves with this modification are reported in Figure˜A12. We observe very similar curve shapes even if the score map is frozen after only a tiny fraction of mask parameters was pruned. This underpins our intuition from Section˜2.2.2 that the pruning path of each parameter is fully determined at very early stage of ACIP.
D.6 Impact of the Score Map – A Trivial One Does Not Work
There are certainly alternative ways to design useful score maps. For example, simply accumulating the gradients of all mask parameters entrywise over an ACIP-run works equally well as the strategy proposed in Section˜2.2.2. It is therefore valid to ask whether one could even design score maps without any optimization. We demonstrate that perhaps the most obvious approach, namely setting the score map equal to the singular values of the weight matrices, does not work very well. Figure˜A13 shows that this training-free approach does not produce any reasonable compressed models and decent performance cannot be easily recovered with LoRA-finetuning. This simple experiment confirms that designing useful score maps is not a trivial endeavour and requires a carefully crafted algorithmic approach.
@@ D.6 Impact of the Score Map – A Trivial One Does Not Work
D.7 Impact of Post-Tuning
Our main experiments are performed with 1k post-tuning steps in ACIP (see the description in Section˜2.2.2 and Appendix˜B). Figure˜A14 shows analogous compression-performance trade-off curves for fewer or no post-tuning steps. We observe that post-tuning can indeed notably increase performance for higher compression ratios.
@@ D.7 Impact of Post-Tuning
D.8 Impact of Individual Layers – Example of LLaMA2-13B
As pointed out in the caption of Table˜A2, the linear layers targeted by ACIP were slightly modified for LLaMA2-13, namely all up projection layers were ignored. Figure˜A15 shows what would happen if they are compressed as well. While the performance predictions for $\geq 0.6$ look decent, the perplexity explodes for stronger compression; note that even additional fine-tuning does not recover a reasonable performance in this situation. We hypothesize that ACIP has pruned one or more singular values of the up projection layers that are crucial for model’s integrity. This finding might be related to the recent work by Yu et al. (2024) on pruning so-called super weights. In any case, ACIP is capable of revealing this undesirable behavior as demonstrated in Figure˜A15.
@@ D.8 Impact of Individual Layers – Example of LLaMA2-13B
D.9 Examples of Score Maps Generated by ACIP
Figure˜A16 and Figure˜A17 show two typical score maps generated by ACIP for LLaMA-7B and Qwen2.5-7B, respectively. A characteristic feature is that attention layers can be pruned more aggressively than the MLP layers. Similarly, we observe non-uniform pruning patterns for layers of the same type across all transformer layers. This confirms the findings of Yuan et al. (2024); Jaiswal et al. (2024) and demonstrates that non-uniform structured compression can be achieved without any feature engineering.
@@ D.9 Examples of Score Maps Generated by ACIP
Figure A16: Example score maps generated by ACIP for LLaMA-7B. The negative values (cf. Algorithm˜1) are normalized to 1-1 for the purpose of visualization.
@@
Figure A17: Example score maps generated by ACIP for Qwen2.5-7B. The negative values (cf. Algorithm˜1) are normalized to 1-1 for the purpose of visualization.
D.10 Examples of Generated Text by Compressed Models
Table˜A6 shows examples of generated text by compressed versions of LLaMA-7B.
@@ D.10 Examples of Generated Text by Compressed Models
70% Apr 04, 2018 at 12:05 pm In the last 5-6 years, Natural Language Systems have been an important direction for my work and the research I do in the course of my PhD. I have been using them to teach myself Swedish and improve my Swedish by machine time in this post. Let’s see if I’ve got it! … Please tell me where is the concept of time in this poem. What is the significance of the poem "I Am Wine" by H.W.C. Carson"? What does he mean by the poem title? What does the stairs, the wine, & the man have in common? What does he mean by "I am wine"? What do you mean by "I am wine"? …
@@
50% Jeffrey Huan, Shu-Hua Ji, Xiao-Hong Li, Xiao-Xiao Li, Layin Ma”s. 2019:5, Springer.0000.0001/13/8/19/12/720-22-F43. Miller A. "The Structure of the Language in Unsupervised Deep-Processing" in Journal of the Acoustical Actors (v. 11(6). 2006:, pp. 278-295). Dong Ming, Sun Jin, He Mei-Chen,’ … Getting to this answer isn’t that hard. The concept of time is all about the study of time. The study of time isn’t so hard to get a hold on. When you’re younger, you can begin thinking of how much you could have had if you had known how much time you had…if only we had it back then …